OpenStudy (anonymous):

f(x)=-2x^2+2x+10, what is the vertex? line of symmetry? if F(1/2)=21/2, minimum or maximum?

OpenStudy (anonymous):

Parabola opens downwards, therefor vertex is a max

OpenStudy (anonymous):

x=-b/2a=-2/2(-2)=1/2 is the x-coordinate of the vertex and the equation for the axis of symmetry

OpenStudy (anonymous):

vertex at (0.5, 10.5)

OpenStudy (anonymous):

line of symmetry: x = 0.5

OpenStudy (anonymous):

thank you.